OpenStudy (anonymous):

two cards are drawn from a pack of cards. is the expected number of hearts drawn larger when the second drawing is made with or without replacement?

OpenStudy (anonymous):

I draw a heart for the first card. Do I have a higher chance of getting another heart if I put that heart back in the deck or not?

OpenStudy (anonymous):

with replacement no.. without then its the same chance

OpenStudy (anonymous):

yes. with replacement gives you a higher number of expected hearts and a higher chance of drawing 2 hearts.

OpenStudy (anonymous):

ok ya I got them mixed up but how do I find the expected number of hearts drawn?

OpenStudy (anonymous):

with replacement: 13/52 x 13/52 or 1/4 x 1/4

OpenStudy (anonymous):

without replacement 13/52 x 12/52

OpenStudy (anonymous):

book says both answers should be 0.5

OpenStudy (anonymous):

i'm sorry...what i gave you is the PROBABILITY of getting 2 hearts

OpenStudy (anonymous):

lol its ok I started to get confused

OpenStudy (anonymous):

ok. with replacement expected number of hearts = expected number on 1st draw + expected number on 2nd draw = 13/52 + 13/52 = 26/52 = 0.5

OpenStudy (anonymous):

without replacement is more complicated 13/52 + 1/4(12/52) + 3/4(13/52)

OpenStudy (anonymous):

why is it 1/4 and 3/4?

OpenStudy (anonymous):

what do each of those number represent because im trying to get a visual

OpenStudy (anonymous):

because the expected number of hearts on the 2nd draw depend on whether or not you got a heart on the first draw. 1/4 th of the time you will get a heart on the first draw: then 12/52 of the remaining cards are hearts 3/4ths of the time you will not get a hear on the first draw: then 13/52 of the remaining cards are hearts.

OpenStudy (anonymous):

hmmm I guess.. thanks for your help

OpenStudy (anonymous):

no problem